Assume that the total revenue of 3 units of a commodity is Birr 24, and

the marginal revenue of the 4th unit is Birr 4. Calculate the average

revenue of 4 units


1
Expert's answer
2021-06-03T13:07:48-0400

Solution:

Average Revenue (AR) = "\frac{TR}{Q}"


Marginal Revenue (MR) = "\frac{\triangle TR}{\triangle Q} or \;TR_{n}-TR _{n-1}"

TR of 3 units = 24

MR of the 4th unit = 4


Derive the Total Revenue (TR) of the 4th unit:

MR of the 4th unit = "TR_{n}-TR _{n-1}"

n = 4

n – 1 = 4 – 1 = 3

MR of the 4th unit = TR4 – TR3

TR3 = 24

4 = TR4 – 24

TR4 = 4 + 24 = 28

Total Revenue of the 4th unit = 28


Therefore,

Average Revenue of the 4 units = "\frac{TR \;of\;4\; units}{4\; units}"


AR of the 4 units = "\frac{28}{4} = 7"


Average Revenue of 4 units = Birr 7